Find the P-value for the indicated hypothesis test. An airline claims that the no-show rate for passengers booked on its flights is less than 5%. Of 380 randomly selected reservations, 18 were no-shows. Find the P-value for a test of the airline's claim. Round to 2 decimals.
Answer #1

H0: p = 0.05

Ha: p < 0.05

Sample proportion = 18 / 380 = 0.0474

Test statistics

z = - p / sqrt( p ( 1 - p) / n)

= 0.0474 - 0.05 / sqrt( 0.05 * 0.95 / 380)

= -0.2326

p-value = P( Z < z)

= P( Z < -0.2326)

= 0.408

p-value = 0.41